OpenStudy (anonymous):

Solve by factoring the trinomial as a binomial square and then use the square root method. Round to the nearest tenth, if necessary. x2 + 16x + 64 = 81 Is it (-17, 1)? I get that but im not sure if its right.

OpenStudy (anonymous):

x^2 + 16 x - 17 = 0 (x - 1)(x + 17) = 0 roots are -17 and +1 as you got

OpenStudy (anonymous):

also ( x + 8)^2 = 81 x + 8 = +- 9 x = -17 or 1